OpenStudy (anonymous):

What is the value of x? 1/5(2x-10)+4x=-3(1/5x-4) a.-2/5 b.-1 1/5 c.-2 d. -5 5/9

OpenStudy (pinklion23):

ill help

OpenStudy (pinklion23):

first do the first side of the equation

OpenStudy (pinklion23):

@josedavid

OpenStudy (pinklion23):

@lvee11

OpenStudy (dtan5457):

\[\frac{ 1 }{ 5 }(2x-10)+4x=-3(\frac{ 1 }{ 5 }x-4)\] Like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2\[4\frac{ 2 }{ 5 }x-2=-\frac{ 3 }{ 5 }x+12\]

OpenStudy (dtan5457):

Yeah, that is correct so far. I'm pretty sure you can solve this yourself lol

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

x=\[-\frac{ 2 }{ 5 }\]

OpenStudy (dtan5457):

your a bit off. once you add -3/5 to the other side. you get 5x. 5x=14

OpenStudy (anonymous):

2 2/5

OpenStudy (dtan5457):

Not exactly..use a calculator. what is 14/5?

OpenStudy (anonymous):

2.8

OpenStudy (dtan5457):

Yes, correct.

OpenStudy (anonymous):

why isn't it in my answer choice

OpenStudy (anonymous):

a.\[-\frac{ 2 }{ 5 }\] b.\[-1 \frac{ 1 }{ 5 }\] c.\[-2\] d. \[-5 \frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

I don't know. I must of typed the question incorrectly but you didn't tell me. For what I wrote above^^ 2.8 is the answer.

OpenStudy (anonymous):

ok thanks